OpenStudy (anonymous):

what is the lcd of 8,3 and 9

OpenStudy (anonymous):

lcd=2^3*3^2=8*9=72

OpenStudy (anonymous):

find a prime factorization of each and then take each distinct factor at the highest power

OpenStudy (anonymous):

8=2^3 3=3^1 9=3^2 the distnct factors are 2 and 3 the highest powers are 3 and 2 repsectively, thus 2^3*3^2=8*9=72

OpenStudy (anonymous):

i did that and that and it told me i was wrong

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=lcd+8%2C3%2C9 check out the above link

OpenStudy (anonymous):

this is the question 3/8+2/3+(-10/9)

OpenStudy (anonymous):

i got -15/72

OpenStudy (anonymous):

oh! we just found the lcd, we didn't add the fractions; now you have to mult each fraction by a useful form of 1 in change each denominator to the lcd:\[\frac{3}{8}*\frac{9}{9}+\frac{2}{3}*\frac{24}{24}-\frac{10}{9}*\frac{8}{8}\]\[=\frac{27+48-80}{72}=\frac{-5}{72}\]this is what i got

OpenStudy (anonymous):

ohhh then i added wrong smh i feel dumb thank you sooo much

OpenStudy (anonymous):

ur welcome